You obtained F1 from a cross of AABB X aabb. F1 produced the following number of gametes

AB: 36 ab: 28 Ab: 14 aB: 18

To test whether A and B gens are linked or not, calculate chi-square value.

Please work out and explain each step!
Answer #1

Answer:

Chi square calculations

Null hypothesis: A and B gens are NOT inked  

Expected values 1:1:1:1 if NOT linked

Genotype

Observed Number

Expected Number

=(Total / sum of ratio)*ratio

χ2 value= (observed- expected)2 / expected

AB

36

(96 / 4 ) * 1= 24

(36-24)2/24= 6

ab

28

(96 / 4 ) * 1= 24

(28-24)2/24= 0.66

Ab

14

(96 / 4 ) * 1= 24

(14-24)2/24= 4.16

aB

18

(96 / 4 ) * 1= 24

(18-24)2/24= 1.5

Total

96

CHI SQUARE VALUE(sum) = 12.32


chi-square value for the observed data= 12.32

Degrees of freedom: (number of groups-1)= (4-1)= 3

Since the chi-square value is not significant the null hypothesis does not hold true

Hence the A and B gens are linked